The proportion of assignments completed within a given day is described by the probability density function f (x) = 12(x - x3) for 0 ≤ x ≤ 1. Find the expected proportion of completed assignments.


1
Expert's answer
2022-03-01T13:40:29-0500

Check


"\displaystyle\int_{-\infin}^{\infin}f(x)dx=\displaystyle\int_{0}^{1}12(x-x^3)dx"

"=[6x^2-3x^4]\begin{matrix}
 1\\
 0
\end{matrix}=6-3=3\not=1"

Let "f (x) = 4(x - x^3)." for "0 \u2264 x \u2264 1."

Check


"\displaystyle\int_{-\infin}^{\infin}f(x)dx=\displaystyle\int_{0}^{1}4(x-x^3)dx"

"=[2x^2-x^4]\begin{matrix}
 1\\
 0
\end{matrix}=2-1=1""E(X)=\displaystyle\int_{-\infin}^{\infin}xf(x)dx=\displaystyle\int_{0}^{1}x(4(x-x^3))dx"

"=[\dfrac{4}{3}x^3-\dfrac{4}{5}x^5]\begin{matrix}
 1\\
 0
\end{matrix}=\dfrac{8}{15}"